On July 1 of the current year, Sammy refinances his home and borrows $240,000. Sammy is required to pay three points on the loan. The loan is secured by the residence and the charging of points is an established business practice in the area. The term of the loan is 30 years, beginning on July 1 of the current year. How much, if any, of the points may Sammy deduct in the current year (rounded)?

$0

$60  

$120

$7,200
Answer #1

Answer : Option C , $120

Explanation;  

$240,000 × 3 % = $7,200 Points Paid

30 Years × 12 Months Per Year = 360 Months in the loan .

$7,200 ÷ 360 = $20 Per Month amortization

This Year 6 Months = $20 × 6 = $120